OpenStudy (anonymous):

How many grams of magnesium metal will react completely with 6.3 liters of 4.5 M HCl? Show all of the work needed to solve this problem. Mg (s) + 2 HCl (aq) yields MgCl2 (aq) + H2 (g)

OpenStudy (anonymous):

Any help would be great. I have no idea what's going on in chemistry.

OpenStudy (nathalyn):

moles HCl = 4.8 x 5 = 24 moles Mg = 24/2 = 12 mass Mg = 12 mol x 24.305 g/mol= 292 g

OpenStudy (anonymous):

Would you be able to describe exactly how to do this?

OpenStudy (nathalyn):

well I wish but no, but what I gave you that's the answer

OpenStudy (anonymous):

Okay. Well, thanks anyway.

OpenStudy (nathalyn):

welcome

OpenStudy (anonymous):

moles= volume(litres) X molarity(Concentration) =6.3 X 4.5 = 28.35 moles mole ratio Mg:HCl =1:2 if 2= 28.35 moles 1= ? \[\frac{ 1\times28.35 }{ 2 }\] Mg has 14.175 moles mass= moles X RAM = 14.175 X23 =326.025g

OpenStudy (anonymous):

different answer? @NathalyN what's wrong with my answer then?